OpenStudy (anonymous):

A cereal manufacturer has a tolerance of .75 ounce for a box of cereal that is supposed to weigh 20 ounces. Write and solve an absolute value inequality that describes the acceptable weights for "20 ounce" boxes. Please note: use '<=' to write a less-than-or-equal-to sign, and '>=' to write a greater-than-or-equal-to sign.

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

it should be \(\Large20-0.75< 20+m<20+0.75\)

OpenStudy (mathmath333):

\(\Large20-0.75\leq 20+m\leq 20+0.75\)

OpenStudy (mathmath333):

simplify it and use \(\Large -a\leq x \leq a\\ \Large x=|a|\)